OpenStudy (anonymous):

what is the square root of y^7

OpenStudy (anonymous):

\[\sqrt {y^7}=\sqrt{y*y^6}=\sqrt{y}\sqrt{y^6}\] can you solve it now?

OpenStudy (anonymous):

what does y represent kim?